# Large Language Model Can Continue Evolving From Mistakes

### v1

@@ Abstract
Large Language Models (LLMs) demonstrate impressive performance in various downstream tasks. However, they may still generate incorrect responses in certain scenarios due to the knowledge deficiencies and the flawed pre-training data. Continual Learning (CL) is a commonly used method to address this issue. Traditional CL is task-oriented, using novel or factually accurate data to retrain LLMs from scratch. However, this method requires more task-related training data and incurs expensive training costs. To address this challenge, we propose the Continue Evolving from Mistakes (CEM) method, inspired by the ’summarize mistakes’ learning skill, to achieve iterative refinement of LLMs. Specifically, the incorrect responses of LLMs indicate knowledge deficiencies related to the questions. Therefore, we collect corpora with these knowledge from multiple data sources and follow it up with iterative supplementary training for continuous, targeted knowledge updating and supplementation. Meanwhile, we developed two strategies to construct supplementary training sets to enhance the LLM’s understanding of the corpus and prevent catastrophic forgetting. We conducted extensive experiments to validate the effectiveness of this CL method. In the best case, our method resulted in a 17.00% improvement in the accuracy of the LLM.
@@
Haokun Zhao, Haixia Han, Jie Shi, Chengyu Du, Jiaqing Liang, Yanghua Xiao Shanghai Key Laboratory of Data Science, School of Computer Science, Fudan University hkzhao23@m.fudan.edu.cn
@@ 1 Introduction
Figure 1: Continual evaluation can help us uncover knowledge deficiency of the LLM and thus evolve the model through continuous learning.
With the exponential expansion of both data and parameter size, large language models (LLMs) demonstrate outstanding performance across a series of downstream tasks Chang et al. (2023). However, LLMs have the potential to generate incorrect outputs, a phenomenon often referred to as ’hallucination’ Huang et al. (2023). In addition, it is important to note that satisfactory answers generated by LLMs cannot be achieved within a single training session Han et al. (2024). Continual refinement and optimization processes are necessary. This optimization process ensures that the LLM can undergo iterative refinement, continuously enhancing its performance.
There are two main reasons for the poor performance of LLM on specific questions. Firstly, some of these inquiries exceed the knowledge boundaries of LLM, indicating either a deficiency in domain expertise or outdated information within the LLM Ren et al. (2023); Zhang et al. (2023). Secondly, the LLM may incorporate erroneous knowledge derived from potentially unreliable sources in the training data, resulting in the generation of inaccuracies or misconceptions that contradict factual information Rae et al. (2021); Zhao et al. (2023).
To rectify the model’s knowledge deficits and ensure alignment with contemporary world knowledge, a common method is periodic training from scratch using the domain or latest corpus, such as the most recently archived Wikipedia. However, this method incurs significant training expenses Patterson et al. (2021). Another method involves pre-training on a corpus that includes accurate information. However, this method may result in a decline in other aspects of the model’s performance Li et al. (2022). This challenge has prompted research into mitigating catastrophic forgetting in continuous learning. Existing studies can be roughly grouped into three categories, e.g., experience replay methods Chaudhry et al. (2019); Wu et al. (2021), regularization-based methods Kirkpatrick et al. (2017a); Chen et al. (2022), and dynamic architecture methods Mallya et al. (2018). Recently, researchers have designed some hybrid methods that take advantage of the aforementioned techniques Chen et al. (2023); He et al. (2024).
In this paper, we focus on the process of uncovering knowledge that the model has not yet mastered, for supplementation. Throughout each iterative phase of model evolution, we identify the limitations of the model’s knowledge and address them in subsequent updates. This idea is inspired by the ’summarizing mistakes’ learning skill observed in human students. An evaluation is conducted initially to identify a student’s weaknesses in a subject. References related to these errors are then provided to assist the student in mastering the associated knowledge. For instance, if a student is struggling with chemistry, they can review questions answered incorrectly in chemistry tests to pinpoint areas needing improvement and extract relevant knowledge, such as the properties of benzene, from the corresponding chapter in the chemistry textbook. This process enhances his performance in chemistry by identifying and learning from weaknesses or areas of unfamiliarity.
In this paper, we propose the Continue Evolving from Mistakes (CEM) method. The aim is to systematically address the model’s knowledge gaps and rectify incorrect knowledge through continuous evolution. Our contributions are summarized as follows:
  • To endow LLMs with the capacity for continuous iterative evolution, we have drawn inspiration from the ’summarize mistakes’ learning skill observed in human students. We introduce CEM, which explores the knowledge boundaries from LLM mistakes, thereby continuously enhancing the model’s capabilities.
  • To assist LLMs in acquiring and comprehending factual knowledge from supplementary corpora, as well as mitigating knowledge forgetting, we propose two strategies for constructing supplementary corpora.
  • We conduct experiments on three open-source LLMs and two complex knowledge Question-Answering (QA) tasks. The results demonstrate the efficacy of the CEM.
@@ 2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examples. Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2021), prompt tuning Li and Liang (2021), and other techniques.
Continual learning with LLM Continual learning (CL) aims to enable a model to learn from a continuous data stream across time while reducing catastrophic forgetting of previously acquired knowledge Biesialska et al. (2020). With CL, a deployed LLM has the potential to adapt to the changing world without costly re-training from scratch Bubeck et al. (2023). Early works Gururangan et al. (2020); Röttger and Pierrehumbert (2021); Lazaridou et al. (2021); Dhingra et al. (2022) empirically analyze continuing LM pre-training on emerging domain or temporal data, showing the potential to update the base LM with new knowledge. Jang et al. (2022b) explicitly categorize world knowledge as timeinvariant, outdated, and new knowledge, which should be retained, acquired, and updated respectively by an LM when learning continually. Jin et al. (2022); Jang et al. (2022a) additionally implement traditional CL methods to alleviate catastrophic forgetting, a phenomenon in which previously learned knowledge or abilities are degraded due to overwritten parameters Kirkpatrick et al. (2017b).
@@ 3 Methods
The objective of this study is to enhance the knowledge base of LLM through continual learning, resulting in iterative improvements in model performance.
@@ 3.1 Task Formulation
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i}$
@@
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplementary corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\}$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and retrieve supplementary corpus.
Subsequently, based on the supplementary corpus $C_{S}$, construct the supplementary training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D)$
where $Construct(\cdot)$ is utilized to build an supplementary training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplementary training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplementary training, thereby achieving continual optimization of the model’s performance.
@@ 3.2 CEM Method
Specifically, the CEM method proposed in this paper involves three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplementary corpora from multiple data sources, and (3) constructing a supplementary training set and fine-tuning the LLM. Fig.2 shows the process.
@@
Collecting Supplementary Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplementary corpora from multiple data sources, including online encyclopedias and search engines.
(1) Search Engine: We submit questions to search engines and use their retrieval techniques to crawl relevant content from the vast corpus of Chinese-language internet. Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We conduct experiments on two search engines, Baidu and Bing, and optimized our corpus crawling strategy.
@@
Specifically, we search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplementary corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. In this experiment, we utilize Bing as the search engine and set $k=4$, indicating that the main textual content from the detailed pages of the top-4 search results with the highest relevance was retrieved. Fig.5 shows process of collecting corpora from search engine.
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. In our experiment, we experiment two online encyclopedias, Baidubaike and Wikipedia.
@@
Specifically, since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplementary corpora.
@@
In our experiment, we use Wikipedia as the online encyclopedia and instruct ChatGPT-3.5 using a few-shot prompt for entity extraction. Appendix Fig.6 shows process of collecting corpora from online encyclopedia.
@@
Constructing Supplementary Training Set and Fine-Tuning LLM. The Supplementary Training Sets are constructed separately based on the Supplementary Corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
@@
After completing one round of supplementary training, the model is retested on the QA dataset to gather a new set of incorrectly answered questions, which serve as the first step for the next round of supplementary training.
@@
It should be noted that supplementary training should not be endless, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement.
3.3 Supplementary Training Data
Corpus Format Definition
@@ 3.3 Supplementary Training Data
The construction of the supplementary training set involves the following four types of corpus.
(1) Format Corpus: Table 1 shows an example of a Format Corpus. Prior to implementing the CEM method for supplementary training, we fine-tune the LLM using Format Corpora. This equips the LLM with the ability to respond to multiple-choice questions in a standard format, facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a certain proportion of Format Corpus during the supplementary training to prevent the LLM’s response skills from declining.
Example 1: A Case of Format Corpus
@@
Table 1: An example of a Format Corpus, designed to illustrate its corpus format.
(2) Supplement Corpus: Corpus obtained through multiple data sources.
(3) Extract Corpus: Table 2 shows an example of an Extract Corpus. Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades. We posit that merely providing the LLM with a supplementary corpus containing knowledge is inadequate. Consequently, we designed the Extract Corpus to train the LLM to effectively capture and comprehend knowledge from supplementary corpora, enhancing its utilization of these materials.
Example 2: A Case of Extract Corpus
@@
Table 2: An example of a Extract Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplementary text has been omitted for the sake of clarity.
(4) General Corpus: Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, we utilize the training data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
@@
as General Corpus to preserve the model’s general capabilities.
Supplementary Training Set Construction Methods. Through the combination of four types of corpora, we have developed three innovative methods for constructing supplementary training sets. These methods are detailed below.
(1) Normal Method: To mitigate the potential catastrophic forgetting resulting from supplementary training, we mix the General Corpus with the Supplement Corpus, ensuring an equal quantity of those two type of corpus across all construction methods,
$D_{N}=Construct(C_{G},C_{S})$
where $Construct(\cdot)$ is utilized for constructing supplementary training set based on corpora, $C_{G}$ represents the General Corpus, $C_{S}$ indicates the Supplement Corpus, and $D_{N}$ denotes the supplementary training set constructed via the Normal Method.
(2) Extract Method: To evaluate whether the Extract Corpus enhances the LLM’s ability to capture and comprehend knowledge from supplementary corpora as expected, and to associate questions with relevant knowledge when answering, we added the Extract Corpus into $D_{N}$ to generate the supplementary training set $D_{E}$.
$D_{E}=Construct(C_{G},C_{S},C_{E})$
We replaced the Extract Corpus with the General Corpus to obtain the control group $D_{E_{C}}$.
$D_{E_{C}}=Construct(C_{G}^{\prime},C_{S})$
(3) Review Method: Under higher temperature parameter, the responses of the LLM tend to be diverse. When conducting repeat QA tests on the LLM, it may change its answers to certain ’less confident’ questions due to the proximity of predicted probabilities. This means that after supplementary training, questions initially answered correctly may be answered incorrectly, resulting in a transition from right to wrong.
Therefore, we strengthen the LLM’s memory of the knowledge implied by correctly answered questions by reviewing the responses provided correctly by the LLM in the previous round.
Specifically, in this round of supplementary training, we sample the correct responses made by the LLM in the previous round and process the correctly answered questions into either ’Format Corpus’ or ’Extract Corpus’ formats. We define it as the Review Corpus, and based on the proportion of ’Format Corpus’ and ’Extract Corpus’ formats in the Review Corpus, we construct $R_{1}$, $R_{2}$, and $R_{3}$,
$C_{R_{1}}=Format(\alpha S_{QA})+Extract((1-\alpha)S_{QA})$
$C_{R_{2}}=Format(S_{QA})$
$C_{R_{3}}=Extract(S_{QA})$
where $\alpha\in(0,1)$, $S_{QA}$ denotes the QAs sampled from the correct responses in the previous round of LLM. $Format(\cdot)$ is employed to process the sampled QAs into Format Corpus format, while $Extract$ is used to process the sampled QAs into Extracted Corpus format.
Note that although Review Corpus looks like a mixture of Extract Corpus and Format Corpus in its compositional form, the corpus of the former comes from the correct responses made by the model and is a proper subset of the test set. In contrast, the corpus used by Extract Corpus and Format Corpus is completely independent of the test set.
We added the Review Corpus into the foundation of $D_{E}$ to generate the supplementary training set $D_{R_{i}}$,
$D_{R_{i}}=Construct(C_{G},C_{S},C_{E},C_{R_{i}}),\quad i=1,2,3$
And we replaced the Review Corpus with the General Corpus to obtain the control group $D_{R_{C}}$.
$D_{R_{C}}=Construct(C_{G}^{\prime\prime},C_{S},C_{E}),$
4 Experiments
4.1 Experimental Settings
Datasets. We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.

### v3

@@ Abstract
As world knowledge evolves and new task paradigms emerge, Large Language Models (LLMs) often fall short of meeting new demands due to knowledge deficiencies and outdated information. Continual Learning (CL) is crucial for keeping LLMs up-to-date and addressing these deficiencies. However, traditional CL approaches struggle to balance task-width generality with task-depth specificity and often lack efficient data collection strategies, leading to increased training costs without addressing the model’s most critical needs.
Inspired by the ‘summarizing mistakes’ learning skill, we propose the Continue Evolving from Mistakes (CEM) method. This iterative approach continually evaluates LLMs to identify knowledge deficiencies based on their mistakes, collecting relevant data from multiple sources to supplement training in a targeted manner. To enhance the model’s utilization of supplemental knowledge and prevent forgetting, we developed three dataset construction strategies that integrate various types of continual pretraining (CPT) data and continual instruction tuning (CIT) data. Extensive experiments demonstrate the efficacy of the CEM method, achieving up to a 17% improvement in LLM accuracy in the best scenarios. Additionally, further experiments confirm the potential of combining CEM with other catastrophic forgetting mitigation strategies, enabling multi-round iterative optimization.
@@
[…]
@@ 1 Introduction
With the exponential growth of data and parameter sizes, coupled with the employment of instruction fine-tuning, Large Language Models (LLMs) have showcased superior performance across a wide array of downstream tasks Chang et al. (2023); Wei et al. (2022); Peng et al. (2023). However, in real-world applications, the ongoing emergence of new knowledge and task requirements necessitates continual task-specific adaptation and knowledge supplementation for LLMs. Otherwise, LLMs are likely to generate incorrect outputs based on outdated or lacking knowledge (i.e., ‘hallucinations’) Huang et al. (2023). Accordingly, Continual Learning(CL) is proposed to ensure LLMs’ alignment with the ever-changing world knowledge.
Given the significant training costs, it is more popular to continually pre-train or fine-tune LLMs on a sequence of emerging domain-specific or task-specific data, rather than periodically training them from scratch on the latest corpora (e.g., the most recently archived Wikipedia) Gururangan et al. (2020b). However, common CL tends to be task-width oriented, whereby the model is trained using new task-relevant data to enhance its applicability across a wider range of tasksGururangan et al. (2020b); Cossu et al. (2022), as opposed to focusing on the continual improvement of the model’s performance within a specific domain or task, i.e., task-depth oriented. In addition, building CL datasets is often challenging. Blind data collection fails to ensure access to the high-quality data that the model most requires for task-solving, and may introduce significant training costs due to irrelevant data.
To address these issues, we propose the Continue Evolving from Mistakes (CEM) method, aiming to provide a data-efficient method for continually enhancing the performance of LLMs on specific tasks through muiltiple evaluation and supplementation with mistake-relevant knowledge. Our method draws inspiration from the ‘summarizing mistakes’ learning skill of human students. For instance, when a student is struggling with math, they can leverage test mistakes to identify lacking knowledge and extract relevant information from the textbook. Through repeated math tests and summaries, improvements in their math scores have been noted. While recent studies have uncovered successful utilization of mistakes in constructing CoT dataAn et al. (2024), thereby enhancing the model’s performance in mathematical reasoning, there exists potential to extend the leveraging of mistakes to diverse tasks and integrate it within CL paradigms.
Recent studies have shown that the poor performance of LLM on specific tasks may result from two factors: (1) issues beyond its knowledge boundaries, indicating a lack of domain expertise Ren et al. (2023); Zhang et al. (2023b), and (2) knowledge errors within LLM, including those stemming from unreliable pre-training data sources and outdated information Rae et al. (2021); Zhao et al. (2023). We argue that mistakes effectively expose the LLM’s knowledge deficiencies and errors. Addressing these directly represents the most efficient approach to enhancing the model. Therefore, at the start of each CEM round, we test the model and gather cases of incorrect or poor performance. We deconstruct these ‘bad cases’ to identify the most lacking or erroneous knowledge points, then obtain supplemental data from search engines and online encyclopedias to address these gaps. To enhance the LLM’s knowledge acquisition from this supplemental data and prevent forgetting, we devised multiple dataset construction strategies utilizing both unsupervised Continual Pre-training (CPT) data and supervised Continual Instruction Tuning (CIT) dataWu et al. (2024). By integrating CEM with strategies to mitigate catastrophic forgetting, we enable iterative, continual model evolution.
Our contributions are summarized as follows: (1) We propose the CEM method. This method explores knowledge boundaries from LLM’s mistakes, endowing LLMs with the capacity for continual iterative evolution. (2) To assist LLMs in acquiring and comprehending factual knowledge from supplemental corpora, and to mitigate knowledge forgetting, we propose three strategies for constructing supplemental training sets. (3) We conduct experiments on multiple open-source LLMs for two complex knowledge Question-Answering (QA) tasks. Across all scenarios, the CEM method consistently exhibit significant improvements, achieving an optimal accuracy enhancement of 17%. (4) We verify that CEM can be used together with methods for mitigating catastrophic forgetting Li et al. (2022), such as replay-based methods Chaudhry et al. (2019); Wu et al. (2021). This ensures that catastrophic forgetting can be suppressed even with multiple iterations.
@@ 2 Related Work
Instruction fine-tuning LMs often struggle to maintain alignment well with human intent after pre-training. However, this issue can be alleviated by employing techniques such as Instruction Fine-Tuning (IFT) Ouyang et al. (2022); Stiennon et al. (2022) and Reinforcement Learning from Human or AI Feedback (RLHF/RLAIF)Gao et al. (2022); Bai et al. (2022). IFT involves training an LM to generate responses based on provided prompts, which might optionally include task instructions. The data used for fine-tuning is usually less than the original pre-training corpus. These datasets used in IFT are often curated through crowd-sourcing or extracted from an LM that is already capable of generating instruction-following examplesHan et al. (2024). Studies have shown that IFT can enhance zero-shot generalization to various unseen tasks. To optimize the fine-tuning process, various strategies have been proposed to make it more parameter-efficient. These strategies include using adapters Hu et al. (2022), prompt tuning Li and Liang (2021), and other techniques.
Continual Pre-training (CPT)
Continual pretraining involves regularly updating LLMs with the latest facts or adapting them to specialized domains. To continually pre-train LLMs to update facts, researchers utilize dynamic datasets to assimilate real-time data from various sources such as news feeds Sun et al. (2019), scholarly articles Cossu et al. (2022), and social media Cossu et al. (2022).
Continual pre-training enhances domain knowledge through two approaches: (1) domain-incremental pre-training, which accumulates knowledge across multiple domains, and (2) domain-specific continual learning, which refines a general model into a domain expert by training on domain-specific datasets and tasks. In domain-incremental pre-training, Cossu et al. (2022) continually pre-trained LLMs on new multimodal data streams, preparing them for various downstream tasks. Ke et al. (2023) introduces a soft-masking mechanism to update LLMs with domain corpora, aiming to boost performance while preserving general knowledge. Notably, there has been research on continuous pretraining for specific domains such as finance Xie et al. (2023) and e-commerce Ma et al. (2023).
Continual Instruction Tuning (CIT)
Continual Instruction Tuning involves continually fine-tuning LLMs to learn how to follow instructions and transfer knowledge for future tasks Zhang et al. (2023a). It can be broadly classified into two categories: 1) task-incremental CIT and 2) domain-incremental CIT.
Task-incremental CIT aims to continually fine-tune LLMs on a series of task-specific instructions to enable them to solve new tasks. A straightforward solution is to continuously generate instruction-tuning data for new tasks and directly fine-tune LLMs on it Wang et al. (2023). However, studies have shown this process would cause a catastrophic forgetting of the learned knowledge and problem-solving skills in previous tasks Kotha et al. (2024).
Domain-incremental CIT aims to continually fine-tuning of LLMs on domain-specific instructions to acquire the knowledge required for solving tasks in new domains. TAPT Gururangan et al. (2020a) adaptively tunes LLMs on domain-specific data from various domains, followed by evaluation of their text classification ability in each domain. AdaptLLM Cheng et al. (2024) enriches the raw training corpus with reading comprehension tasks, helping the model leverage domain-specific knowledge while enhancing prompting performance.
@@ 3 Methods
The objective of this work is to enhance the knowledge of LLM through continual learning, resulting in iterative improvements in model performance.
@@ 3.1 Task Formulation
$\widetilde{r_{i}}=\mathcal{M}_{0}(q_{i}),\quad Ans(\widetilde{r_{i}})\neq a_{i},$
@@
Then, for each collected question $q_{i}$ that was answered incorrectly by the LLM, collect supplemental corpus $C_{S_{i}}$,
$C_{S_{i}}=Search(q_{i}),\quad C_{S}=\{C_{S_{1}},C_{S_{2}},C_{S_{3}},\dots\},$
where $Search(\cdot)$ is used to search for relevant knowledge based on the question and obtain supplemental corpus.
Subsequently, based on the supplemental corpus $C_{S}$, construct the supplemental training set $D$ to fine-tune model $\mathcal{M}_{0}$,
$D=Construct(C_{S}),\quad\mathcal{M}_{1}=Train(\mathcal{M}_{0},D),$
where $Construct(\cdot)$ is utilized to build an supplemental training set, $Train(\cdot)$ denotes fine-tuning. $\mathcal{M}_{1}$, represents the model that has undergone one iteration of supplemental training.
We can continue to employ $\mathcal{M}_{1}$ as the reasoning model on the test set $\mathcal{Q}$, for iterative supplemental training, thereby achieving continual optimization of the model’s performance.
@@ 3.2 CEM Method
We propose the CEM method, involving three main steps: (1) testing the LLM and collecting incorrectly answered questions, (2) collecting supplemental corpora from multiple data sources, and (3) constructing a supplemental training set and fine-tuning the LLM. Figure 1 shows the process.
@@
Collecting Supplemental Corpora from Multiple Data Sources. We argue that the questions answered incorrectly by the LLM reveal knowledge deficiencies. Based on the implicit information contained within questions incorrectly answered by the LLM, we collect supplemental corpora from multiple data sources, including search engines and online encyclopedias.
@@
(1) Search Engine: Search engines are characterized by fast response time, high relevance matching, and diverse and rich content. We search for questions that the LLM answers incorrectly. The search engine ranks results based on relevance. We select the top-k results with the highest relevance and access their detailed pages to crawl the main textual content as supplemental corpora. To prevent the LLM from simply memorizing answers, we filter out search results that may contain both the original question and its answer. Additionally, we optimize the strategy for crawling the corpus to filter out irrelevant information and spam, such as advertisements.
In subsequent experiments, we utilize Bing as the search engine and set $k=4$. Figure 4 in Appendix shows a case of corpus collection from Bing.
@@
(2) Online Encyclopedia: Compared to using search engines for retrieval, online encyclopedias provide more credible and professional search results, despite having lower hit rates than search engines. Since the organization of encyclopedias typically relies on keyword entries, we conduct entity extraction from the questions and options of the collected incorrectly answered questions, using the extracted entity keywords for retrieval. The main text of the retrieved results serve as supplemental corpora.
@@
In subsequent experiment, we use Wikipedia as the online encyclopedia and instruct GPT-3.5 using a few-shot prompt for entity extraction. Figure 5 in Appendix shows a case of corpus collection from Wikipedia.
@@
Constructing Supplemental Training Set and Fine-Tuning LLM. The supplemental training sets are constructed separately based on the supplemental corpus collected from Wikipedia and Bing. They are then employed to fine-tune the LLM. The construction method of the training set will be explained in detail in Section 3.3.
@@
After completing one round of supplemental training, the model is retested on the QA dataset to evaluate its performance and gather a new set of incorrectly answered questions. This can serve as the first step for the next round of supplemental training.
@@
It is important to note that supplemental training should not be never-ending, as this may lead to severe catastrophic forgetting and the boundary effect of diminishing accuracy improvement. In the following experiments, we will examine how the number of supplemental training rounds affects model improvement.
3.3 Supplemental Training Data
@@
Definition of Corpus Type. The construction of the supplemental training set involves the following five types of corpus.
(1) Supplemental Corpus($C_{\underline{S}}$): Supplemental Corpus obtained from multiple data sources contains background knowledge that the LLM lacks and is used similarly to continuous pre-training Gupta et al. (2023).
Example 1: A Case of Normative Corpus
@@
Table 1: A case of a Normative Corpus, designed to illustrate its corpus format.
Example 2: A Case of Extractive Corpus
@@
Table 2: A case of a Extractive Corpus, designed to illustrate its corpus format. The ellipsis indicates that some supplemental text has been omitted for clarity.
(2) Normative Corpus($C_{\underline{N}}$): To equip the LLM with the ability to respond to multiple-choice questions in a normative format, we sample multiple-choice questions that are independent of the test set and process them into the format shown in Table 1 to build Normative Corpus. Prior to supplemental training, we finetune the LLM using Normative Corpus(i.e. pre-finetuning), facilitating the extraction of answers from the LLM’s responses and allowing for the collection of incorrectly answered questions. To maintain this ability, we incorporate a small proportion of Normative Corpus during the supplemental training to prevent the LLM’s response skills from declining.
(3) Extractive Corpus($C_{\underline{E}}$): Inspired by the human learning process: despite studying from the same textbooks, students who are better at extracting knowledge from these materials always achieve superior grades, we believe that training models to effectively utilize materials is as crucial as acquiring them.
To improve the LLM’s ability to capture and comprehend knowledge from the Supplemental Corpus, we sample multiple-choice questions that are independent of the test set and the Normative Corpus. These questions are processed into the format shown in Table 2 to build the Extractive Corpus.
(4) Review Corpus($C_{\underline{R}}$): The responses of the LLM tend to be diverse under high temperature Caccia et al. (2018). During repeated QA testing, LLM may change its answers to certain ‘less confident’ questions due to the proximity of predicted probabilities. Consequently, after supplemental training, questions initially answered correctly may be answered incorrectly, thus affecting the accuracy.
Therefore, to strengthen the LLM’s memory of the knowledge implied by correctly answered questions, we sample the responses provided correctly of the LLM in the previous round and process some of them in Normative Corpus format and the rest in Extractive Corpus format to build the Review Corpus.
The composition of the Review Corpus can be described as:
$C_{R}=Normative(\alpha r)+Extractive((1-\alpha)r),$
where $\alpha\in[0,1]$, denotes the proportion of responses processed into the Normative Corpus format. $r$ denotes the correct responses in the previous round of LLM. $Normative(\cdot)$ converts responses into the Normative Corpus format, and $Extractive(\cdot)$ converts responses into the Extractive Corpus format.
(5) General Corpus($C_{\underline{G}}$): Replaying previous training data can effectively mitigate catastrophic forgetting when training LLM on new dataHe et al. (2021). Therefore, to preserve the model’s general capabilities on other tasks, we use the data from ShareGPT dataset  [^1] https://huggingface.co/datasets/Aeala/ShareGPT_Vicuna_
@@
as General Corpus.
Supplemental Training Set Construction Strategy. We have developed three innovative strategies for constructing supplemental training sets by combining five types of corpus. The strategies are detailed below.
(1) Plain Strategy: To mitigate the potential catastrophic forgetting and to maintain the LLM’s ability to provide normative responses, we mix three types of corpus to construct the supplemental training set:
$D_{P}=Concat(C_{S},C_{N},C_{G}),$
where $C_{S}$ denotes the Supplemental Corpus, $C_{N}$ denotes the Normative Corpus, $C_{G}$ denotes the General Corpus, $Concat{(\cdot)}$ merges various types of corpus together and $D_{P}$ denotes the supplemental training set constructed via the Plain Strategy.
(2) Extractive Strategy: To enhance the LLM’s ability to capture and comprehend knowledge from supplemental material, we integrate the Extractive Corpus:
$D_{E}=Concat(C_{S},C_{N},C_{G},C_{E}),$
where $C_{E}$ denotes the Extractive Corpus and $D_{E}$ denotes the supplemental training set constructed via the Extractive Strategy.
(3) Review Strategy: To enhance the model’s retention of acquired knowledge and prevent correct responses from being altered to incorrect ones over multiple iterations, we integrate the Review Corpus:
$D_{R}=Concat(C_{S},C_{N},C_{G},C_{E},C_{R}),$
where $C_{R}$ denotes the Review Corpus and $D_{R}$ denotes the supplemental training set constructed via the Review Strategy.
4 Experimental Setup
4.1 Datasets
We conduct experiments on two question-answering datasets, including Xiezhi [^2] https://github.com/mikegu721/xiezhibenchmark and CMMLU [^3] https://github.com/haonan-li/CMMLU.